joan's brother is three times as old as she. in 5 years his age will be twice hers
Answer by TimothyLamb(4379)

x = joan age now
y = bother age now
---
y = 3x
y + 5 = 2(x + 5)
---
put the system of linear equations into standard form
---
3x - y = 0
2x - y = -5
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 5
y = 15
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---